# Validation of the COSINE-100U NaI(Tl) Encapsulation for Low-Temperature Operation in Liquid Scintillator

### v1

@@ 2 Experimental setup
For the optical interface, 2 mm thick silicone pads (Eljen EJ-560) were coupled to the top and bottom faces of the crystal. Two high-quantum-efficiency 3-inch PMTs (Hamamatsu R12669SEL) were attached to these pads to maximize light collection. The curved side of the crystal was wrapped in several layers of soft, Polytetrafluoroethylene (PTFE)-based Tetratex, which serves as a diffuse reflector to contain scintillation light. A PTFE holder secures the mechanical coupling between the crystal and PMTs, ensuring uniform optical contact. The entire assembly was centered in a copper support and enclosed in a copper housing matching the COSINE-100U design [23]. Crucially, the housing is sealed using Viton O-rings selected for their chemical compatibility with LS and mechanical resilience at $30^{\circ}\text{C}$. The overall assembly procedure is illustrated in Fig. 1.
@@ 3.1.1 Decay time measurement
The extracted decay constants and component fractions are summarized in Table 1. The fast decay constant ($\tau_{1}$) remains consistent at approximately 200 ns at both temperatures. The intermediate decay constant ($\tau_{2}$) is also stable, with a value on the order of 1000 ns. However, the low-temperature waveform exhibits an additional slow component ($\tau_{3}$) on the order of 5000 ns. Notably, the contribution of the fast component decreases dramatically from 89% at RT to 29% at $33^{\circ}\text{C}$.
Table 1: Fitted decay constants and component fractions ($f_{1}$ and $f_{2}$) for the multi-exponential model applied to accumulated $^{241}\text{Am}$ 59.54 keV waveforms at room temperature and $33^{\circ}\text{C}$. At room temperature, a two-exponential model provides an adequate description, so no $\tau_{3}$ parameter is included. At $33^{\circ}\text{C}$, a three-exponential model is used. Quoted uncertainties are statistical and rounded to one decimal place.
Temperature $\tau_{1}$ [ns] $\tau_{2}$ [ns] $\tau_{3}$ [ns] $f_{1}$ [%] $f_{2}$ [%]
Room temp. $220.4\pm 0.6$ $1030.0\pm 4.3$ - $88.8\pm 0.1$ $11.2\pm 0.1$
$33^{\circ}$C $208.9\pm 2.4$ $1244.1\pm 4.7$ $5048.8\pm 29.6$ $29.4\pm 0.4$ $54.5\pm 0.2$

### v2

@@ 2 Experimental setup
For the optical interface, 2 mm thick silicone pads (Eljen EJ-560) were coupled to the top and bottom faces of the crystal. Two high-quantum-efficiency 3-inch PMTs (Hamamatsu R12669SEL) were attached to these pads to maximize light collection. The curved side of the crystal was wrapped in several layers of soft, Polytetrafluoroethylene (PTFE)-based Tetratex, which serves as a diffuse reflector to contain scintillation light. A PTFE holder secures the mechanical coupling between the crystal and PMTs, ensuring uniform optical contact. The entire assembly was centered in a copper support and enclosed in a copper housing matching the COSINE-100U design [23]. Crucially, the housing is sealed using Viton O-rings selected for their chemical compatibility with LS and mechanical resilience at $30^{\circ}\text{C}$. The overall assembly procedure is illustrated in Fig. 1. Figure 2 shows a schematic of the completed copper encapsulation with the main dimensions indicated in units of mm. The copper housing is sealed using O-rings at both end-cap (lid–body) interfaces to ensure an airtight environment and prevent the ingress of ambient humidity.
@@
Figure 2: Schematic cross-sectional view of the fully encapsulated NaI-037 detector geometry. The NaI(Tl) crystal (green) is optically coupled to two 3-inch PMTs (magenta), with the PMT bases shown in gray. O-rings (black) at both lid–body interfaces provide the seal between the copper lids and the copper body. A PTFE (Teflon) support structure holds the crystal–PMT assembly inside the upgraded COSINE-100U copper housing; the main encapsulation dimensions are indicated in units of mm.
@@ 3.1.1 Decay time measurement
The extracted decay constants and component fractions are summarized in Table 1. The component fractions are defined from the fitted exponential components using the relative areas of each term: $f_{i}=100\times A_{i}\tau_{i}/(A_{1}\tau_{1}+A_{2}\tau_{2})$ for the RT two-exponential model in Eq. (3.1), and $f_{i}=100\times B_{i}\tau_{i}/(B_{1}\tau_{1}+B_{2}\tau_{2}+B_{3}\tau_{3})$ for the LT three-exponential model in Eq. (3.2); therefore, $f_{1}+f_{2}=100\%$ (RT) and $f_{1}+f_{2}+f_{3}=100\%$ (LT). The fast decay constant ($\tau_{1}$) remains consistent at approximately 200 ns at both temperatures. The intermediate decay constant ($\tau_{2}$) is also stable, with a value on the order of 1000 ns. However, the low-temperature waveform exhibits an additional slow component ($\tau_{3}$) on the order of 5000 ns. Notably, the contribution of the fast component decreases dramatically from 89% at RT to 29% at $33^{\circ}\text{C}$.
Table 1: Fitted decay constants and component fractions for the multi-exponential model applied to accumulated $^{241}\text{Am}$ 59.54 keV waveforms at room temperature and $33^{\circ}\text{C}$. At room temperature, a two-exponential model is used, so $f_{3}$ is not defined. Quoted uncertainties are statistical and rounded to one decimal place.
Temperature $\tau_{1}$ [ns] $\tau_{2}$ [ns] $\tau_{3}$ [ns] $f_{1}$ [%] $f_{2}$ [%] $f_{3}$ [%]
Room temp. $220.4\pm 0.6$ $1030.0\pm 4.3$ - $88.8\pm 0.1$ $11.2\pm 0.1$ -
$33^{\circ}$C $208.9\pm 2.4$ $1244.1\pm 4.7$ $5048.8\pm 29.6$ $29.4\pm 0.4$ $54.5\pm 0.2$ $16.1\pm 0.4$